# HYDRODYNAMIC SIMULATIONS OF TIDAL DISRUPTION ENCORES

### v1

@@ I Introduction
Nuclear star clusters (NSCs), made up of very dense concentrations of stars and the compact objects they leave behind, are often found in the central regions of galaxies surrounding the central massive black hole (MBH). In particular, the fraction of nucleated galaxies is found to be especially large, $\sim 90\%$, in early-type galaxies with masses $\sim 10^{9}\,\mathrm{M}_{\odot}$, albeit dropping at both the high and low end of the mass distribution [1]. The star density in NSCs is very high, possibly reaching $\sim 10^{6}\,\mathrm{M}_{\odot}$ pc-3 at a radial distance of $\sim 0.1$ pc from the core and following a very steep profile, $\propto r^{-1}-r^{-3}$ (see e.g. Neumayer et al. 2 for a comprehensive review). For our Galactic NSC, the stellar density is observed to be $\sim 10^{6}\,\mathrm{M}_{\odot}$ pc-3 even at a distance of 0.25 pc [3].
@@
Ryu et al. [13] studied the evolution of the TDEE analytically, assuming ballistic orbits for the stellar debris. They found that the temporal evolution of the fallback rate of the debris can vary between the standard $t^{-5/3}$ power law to much steeper decays. If the first flare from the micro-TDE is missed, then the TDEE can show up as an ’impostor’ among the more typical TDEs from MBHs. The variety of fallback rates exhibited by these encores may indeed provide a natural explanation for a wide range of decay slopes $t^{-1}-t^{-4}$; 14] of observed light curves of optical/UV TDE candidates which are in tension with the predictions of the classical TDE model [e.g., 15].
@@ II Analytic Expectations - Direct Encore or Ring?
Figure 2: Rings are likely to form to the left of the dashed vertical lines, where $|\epsilon_{\mathrm{sBH}}|>\Delta\epsilon$. To the right of the lines, direct path encores are possible depending on the specific orbital configuration of the three bodies.
@@ III.1 Stellar model
We created a stellar model assuming near-solar metallicity ($Z=0.02$) using the stellar evolution code MESA [version r24.03.1; 20, 21, 22]. We followed the prescriptions for overshoot, convection, semiconvection, and thermohaline mixing from [23], and employed the Ledoux criterion [24] to identify the convective region boundary. We evolved a main-sequence star with a mass of $1\,\mathrm{M}_{\odot}$ up to when the central hydrogen mass fraction reached 0.3. At that time, the star’s surface radius is $1\,\mathrm{R}_{\odot}$ as anticipated. We show the density and Hydrogen mass fraction profiles of the stellar model in Appendix B.
@@ III.3 Bolometric Luminosity
where $H$ is the vertical density-weighted scale height at the given $r$ and $\phi$. The diffusion time represents an “average” timescale across the entire vertical column, in the sense that it does not depend on height. While this is admittedly an approximate estimate, it is robust against uncertainties in the exact photosphere location, which is a limitation inherent to estimates based on adiabatic hydrodynamical simulations.
@@ IV Direct Encore Results
We report the highest encore luminosity of $\approx 10^{41}$ erg/s, corresponding to roughly the Eddington limit for a $10^{3}M_{\odot}$ MBH, and the qualitatively most complex debris trajectories for $\pi/2\leq\theta\leq\pi$. These TDEEs correspond to disruptions occurring closer to the MBH. Although our angle $\theta$ has a slightly different definition than that of Ryu et al. [13], we confirm the most dramatic plunging angle for the debris to be $\theta={3\pi}/{4}$.
Because our version of AREPO does not include radiation transfer, the late time luminosity decay of our TDEE cannot be well predicted within simulations. However, we note that our simulation reach a rotationally stable disk by the end of an sBH orbital period. This final state motivates an understanding of late timescale TDEEs as being fuelled via viscous transfer of angular momentum.
In Fig. 6, we show the viscous timescale distribution for the encore debris as a proxy for the time they would reach the MBH. Accretion of this debris to the MBH can yield an additional source of luminosity on the freefall timescale. As an order of magnitude estimate, we can write $L_{\rm acc}\sim 5\times 10^{44}{\rm erg~s^{-1}}(\eta/0.1)(M_{\rm debris}/0.1M_{\odot})(t_{\rm visc}/1{\rm yr})^{-1}$, where $\eta$ is an accretion efficiency factor. Accretion of the debris onto the MBH, depending on the efficiency, can therefore result in an hyper-Eddington accretion flow for $M_{\rm MBH}=10^{3}\,\mathrm{M}_{\odot}$, possibly driving an outflow/jet, and yield a rather brilliant encore flare.
Figure 6: Distribution of viscous timescales for the debris that reach the MBH in some of the direct encore cases. The dashed lines, sharing the same colors, indicate the range within which 95% of the debris mass is enclosed.
@@ V Ring Encore Results
Figure 9: Mass distribution in the TDEE rings, evaluated at $t\approx 5\mathcal{P}$. The radial extent of the debris containing $95\%$ of the mass is $\sim 26r_{\mathrm{g}}$ for the $10^{-5}$ pc ring, and $\sim 2.5\times 10^{3}r_{g}$ for the ring centred at $10^{-4}$ pc.
@@ V.1 Width and Timescale of the Ring
Figure 10: Viscous and diffusive timescales calculated at $t\approx 5\mathcal{P}$ for both ring simulations. Shaded regions enclose 95% of the debris mass, highlighting the areas of greatest significance. $t_{\mathrm{diff}}$ peaks at the tori centers, corresponding to the densest regions.
@@ VI.1 Observational Implications
For direct encores, our simulations show that debris stream self-intersections can occur as early as one week after the micro-TDE, scaling as $(R_{\rm bin}/10^{-5}\,\mathrm{pc})^{3/2}(M_{\rm{MBH}}/10^{3}\,\mathrm{M}_{\odot})^{-1/2}$. This stage in the light curve corresponds to a sharp luminosity rise from shocks, observable in the optical/UV band (see Fig. 5). While the folding of debris streams into self-intersections near the MBH is complex and highly sensitive to initial conditions, the peak luminosity remains robust against these variations.

### v2

@@ I Introduction
Nuclear star clusters (NSCs), made up of very dense concentrations of stars and the compact objects they leave behind, are often found in the central regions of galaxies surrounding the central massive black hole (MBH). In particular, the fraction of nucleated galaxies is found to be especially large, $\sim 90\%$, in early-type galaxies with masses $\sim 10^{9}\,\mathrm{M}_{\odot}$, albeit dropping at both the high and low end of the mass distribution [29]. The star density in NSCs is very high, possibly reaching $\sim 10^{6}\,\mathrm{M}_{\odot}$ pc-3 at a radial distance of $\sim 0.1$ pc from the core and following a very steep profile, $\propto r^{-1}-r^{-3}$ (see e.g. Neumayer et al. 15 for a comprehensive review). For our Galactic NSC, the stellar density is observed to be $\sim 10^{6}\,\mathrm{M}_{\odot}$ pc-3 even at a distance of 0.25 pc [4].
@@
Ryu et al. [25] studied the evolution of the TDEE analytically, assuming ballistic orbits for the stellar debris. They found that the temporal evolution of the fallback rate of the debris can vary between the standard $t^{-5/3}$ power law to much steeper decays. If the first flare from the micro-TDE is missed, then the TDEE can show up as an ’impostor’ among the more typical TDEs from MBHs. The variety of fallback rates exhibited by these encores may indeed provide a natural explanation for a wide range of decay slopes $t^{-1}-t^{-4}$; 6] of observed light curves of optical/UV TDE candidates which are in tension with the predictions of the classical TDE model [e.g., 5].
@@ II Analytic Expectations - Direct Encore or Ring?
Figure 2: The angle $\varphi$ subtended by the MBH’s ISCO versus the radial separation $R_{\mathrm{bin}}$ for a variety of MBH masses. Rings are likely to form to the left of the dashed vertical lines, where $|\epsilon_{\mathrm{sBH}}|>\Delta\epsilon$. To the right of the lines, direct path encores are possible depending on the specific orbital configuration of the three bodies.
@@ III.1 Stellar model
We created a stellar model assuming near-solar metallicity ($Z=0.02$) using the stellar evolution code MESA [version r24.03.1; 18, 19, 11]. We followed the prescriptions for overshoot, convection, semiconvection, and thermohaline mixing from [2], and employed the Ledoux criterion [13] to identify the convective region boundary. We evolved a main-sequence star with a mass of $1\,\mathrm{M}_{\odot}$ up to when the central hydrogen mass fraction reached 0.3. At that time, the star’s surface radius is $1\,\mathrm{R}_{\odot}$ as anticipated. We show the density and Hydrogen mass fraction profiles of the stellar model in Appendix B.
@@ III.3 Bolometric Luminosity
where $H$ is the vertical density-weighted scale height at the given $r$ and $\phi$. The diffusion time represents an “average” timescale across the entire vertical column, in the sense that it does not depend on height. While this is admittedly an approximate estimate, it is robust against uncertainties in the exact photosphere location, which is a limitation inherent to estimates based on adiabatic hydrodynamical simulations. In Appendix X, we then expand on the results presented here and use our best estimate of the photospheric location in order to compute the spectral energy distributions associated with different TDEE morphologies.
@@ IV Direct Encore Results
We report the highest encore luminosity of $\approx 10^{41}$ erg/s, corresponding to roughly the Eddington limit for a $10^{3}M_{\odot}$ MBH, and the qualitatively most complex debris trajectories for $\pi/2\leq\theta\leq\pi$. These TDEEs correspond to disruptions occurring closer to the MBH. Although our angle $\theta$ has a slightly different definition than that of Ryu et al. [25], we confirm the most dramatic plunging angle for the debris to be $\theta={3\pi}/{4}$. We remark that, because our version of AREPO does not include radiation transfer, the late time luminosity change of our TDEE cannot be accurately predicted with our simulations. We restrict radiative losses - estimated via the time integrated luminosity - to $\lesssim 5\%$ of the sytem’s total thermal energy, to ensure that the hydrodynamical evolution remains robust regardless of whether radiative cooling is explicitly included.
We note that our simulation reaches a rotationally stable disk by the end of an sBH orbital period. This final state motivates an understanding of late timescale TDEEs as being fuelled via viscous transfer of angular momentum. In Fig. 6, we show the viscous timescale distribution for the encore debris as a proxy for the time they would reach the MBH. Accretion of this debris to the MBH can yield an additional source of luminosity on the freefall timescale. As an order of magnitude estimate, we can write $L_{\rm acc}\sim 5\times 10^{44}{\rm erg~s^{-1}}(\eta/0.1)(M_{\rm debris}/0.1M_{\odot})(t_{\rm visc}/1{\rm yr})^{-1}$, where $\eta$ is an accretion efficiency factor. Accretion of the debris onto the MBH, depending on the efficiency, can therefore result in an hyper-Eddington accretion flow for $M_{\rm MBH}=10^{3}\,\mathrm{M}_{\odot}$, possibly driving an outflow/jet, and yield a rather brilliant encore flare.
Figure 6: Distribution of viscous timescales (Eq. 7) with respect to the debris location $r$ for a subset of direct encore simulations. The dashed lines, sharing the same colors, indicate the range within which 95% of the debris mass is enclosed. We may anticipate a moderate enhancement of the accretion rate onto the MBH within the approximate timescale of the optical/UV encore flare.
@@ V Ring Encore Results
Figure 9: Cumulative radial mass distribution of TDEE rings, evaluated after the torus formation at $t\approx 5\mathcal{P}$. The shaded regions indicate the radial extent of the debris enclosing $95\%$ of the debris mass, corresponding to $\sim 26r_{\mathrm{g}}$ for the $10^{-5}$ pc ring and $\sim 2.5\times 10^{3}r_{g}$ for the ring centered at $10^{-4}$ pc.
@@ V.1 Width and Timescale of the Ring
Figure 10: Diffusive cooling and viscous timescales (Eq. 6 and Eq. 7 respectively) as a function of the radial distance at $t\approx 5\mathcal{P}$ for both ring simulations. Shaded regions enclose 95% of the debris mass, highlighting the areas of greatest significance. $t_{\mathrm{diff}}$ peaks at the tori centers, corresponding to the densest regions.
@@ VI.1 Observational Implications
For direct encores, our simulations show that debris stream self-intersections can occur as early as one week after the micro-TDE, scaling as $(R_{\rm bin}/10^{-5}\,\mathrm{pc})^{3/2}(M_{\rm{MBH}}/10^{3}\,\mathrm{M}_{\odot})^{-1/2}$. This stage in the light curve corresponds to a sharp luminosity rise from shocks, observable in the optical/UV band (see Fig. 5 in Appendix X). While the folding of debris streams into self-intersections near the MBH is complex and highly sensitive to initial conditions, the peak luminosity remains robust against these variations.
@@ IX Appendix B: MESA Stellar Model
X Appendix C: Spectral Properties of TDEEs
Using the upward flux $F_{\mathrm{rad}}(r,\phi)$ described in section III.3, together with the temperature at the photosphere along the virtual grid $T_{\mathrm{ph}}(r,\phi)$, we can construct approximated spectral energy distributions (SEDs) associated with different TDEE morphologies.
We discretize the photosphere into surface elements indexed by $k$, each characterized by a photospheric temperature $T_{\mathrm{ph},k}$ and an associated bolometric luminosity contribution $L_{k}$, obtained by integrating the upward radiative flux over the corresponding area element,
$L_{k}\equiv\int_{\Delta A_{k}}F_{\mathrm{rad}}(r,\phi)\,dA.$ (13)
To construct a continuous approximation to the emergent spectrum, we bin the photospheric temperatures into a set of 500 temperature bins $\{T_{i}\}$ with centers $T_{i}$. The luminosity associated with each temperature bin is computed as
$L(T_{i})=\sum_{k\in i}L_{k}\,$, (14)
where the sum runs over all surface elements $k$ whose photospheric temperatures fall within the $i$-th temperature bin.
Each temperature bin is assumed to emit as a blackbody at temperature $T_{i}$. The specific luminosity contributed by a given bin is therefore
$L_{\nu,i}=L(T_{i})\,\frac{B_{\nu}(T_{i})}{\sigma T_{i}^{4}},$ (15)
where $B_{\nu}(T)$ is the Planck function and $\sigma$ is the Stefan–Boltzmann constant.
Although we normalize by $\sigma T_{i}^{4}$ so that integrating $L_{\nu,i}$ over frequency should recover the bolometric luminosity $L(T_{i})$, we find that the integral of our SEDs does not exactly retrieve the bolometric luminosities in Section III.3. This discrepancy is due to variations in the calculation of $L(t)$ from the radiative energy contained in a column of the virtual grid (4) versus a calculation of $L(t)$ from solely the photosphere temperature. For the results provided in the main text, we avoided calculating the bolometric luminosity from $T_{\mathrm{ph}}$ to mitigate AREPO ’s incomplete radiation transfer capabilities, but this agnostic approach is not possible for spectral analysis. Thus, we present these findings with caution since our simulations do not include radiation transfer and the virtual grid does not resolve the photosphere with high precision; however, we believe these estimates can still be useful to broadly categorize the TDEE spectral observability.
The total spectral energy distribution is then obtained by summing over all temperature bins,
$L_{\nu}=\sum_{i}L(T_{i})\,\frac{B_{\nu}(T_{i})}{\sigma T_{i}^{4}}.$ (16)
This procedure is repeated for several simulation times, to construct a sequence of SEDs that illustrate the spectral evolution of different TDEE morphologies (Figures 14-16).
In Figure 14, we show the SED results for our fiducial choice of direct encore for the case $\theta=3\pi/4$. Before the start of the encore, our simulations indicate a shift to lower frequencies and luminosities from the initial micro TDE. As the encore begins around $t=1.7\;\rm{days}$, the spectral peak broadens toward higher frequencies and subsequently increases in luminosity. As time progresses beyond the onset of the encore, the peak luminosity decreases again. This spectral shift is explained by encore shocks causing a secondary peak at $\nu\sim 5\times 10^{15}\;\mathrm{Hz}$ whose luminosity increases between $t=1.7$ and $t=7.3$ days as the shocks persist. As this encore peak grows, the initial peak from the micro TDE continues to cool into the optical band, so that by $t=7.3$ the encore peak becomes dominant. For $t>7.3$ days the debris undergoes adiabatic cooling, although specific dynamic idiosyncrasies may imprint minor features on the subsequent evolution of the SED.
Figure 14: SED for the direct encore case of Figure 4, showed at five different times in order to illustrate the temporal evolution of the SED. Colored regions indicate approximate boundaries for optical (yellow), UV (blue), and X-ray (red) bands.
For other direct encore configurations, we note that the spectral evolution follows a similar trend of peak widening toward higher frequencies and then increasing peak luminosity; however, specific hydrodynamic effects can imprint observable differences on the SED evolution.
In the ring encore scenario, the temporal evolution of the SED is comparatively simpler. In Figure 15, we show SEDs for the $R_{\mathrm{bin}}=10^{-5}\,\mathrm{pc}$ ring encore. For this morphology, encore shocks shift the peak toward higher frequencies. As the shocks associated with disk/torus formation continue after approximately one orbital period, adiabatic expansion cools the debris culminating in the SED peak shifting toward lower frequencies.
Figure 15: SED corresponding to the temporal evolution of the encore of Figure 7. Plotting conventions are the same as in Figure 14; however, because this case corresponds to a ring encore with orbital period $\mathcal{P}\approx 1.1\mathrm{\;days}$, the spectral evolution is markedly different.
When encores happen in less gravitationally bound cases like the $R_{\mathrm{bin}}=10^{-4}\,\mathrm{pc}$ ring encore shown in Figure 16, cooling can already be significant before encore onset, similarly to the direct encore case. This makes the encore more pronounced, as the spectral peak luminosity exhibits larger fluctuations on month-long timescales.
Figure 16: SED corresponding to the temporal evolution of the encore of Figure 8. The SED trends are similar to those of Figure 16, but the longer orbital period ($\mathcal{P}\approx 34\mathrm{\;days}$) allows for more cooling before the onset of shocks in the encore.
While these SEDs are order-of-magnitude estimates, our results highlight TDEEs as compelling targets for multi-wavelength astronomy, as distinct structures can produce spectral peaks in the optical, UV, and X-rays. Moreover, temporal variability in the spectral emission may provide a novel avenue for probing the astrodynamics and population properties of nuclear star clusters.